# booktabs, siunitx

### v1

@@ 3 Generation of scalar-induced secondary GWs and the PTA data
As is well known, primordial GWs are the transverse and traceless components of the tensor perturbations in the metric describing the FLRW universe. There can be different origins for GWs in the early universe. At the linear order in perturbation theory, the scalar and tensor perturbations evolve independently. However, the tensor perturbations at the second order can be sourced by the scalar perturbations at the first order. As a result, whenever the amplitudes of the scalar perturbations are enhanced on small scales so that the PBHs are formed abundantly, they induce tensor perturbations of significant strengths at the second order, often leading to a peak in the spectral density of GWs (see, for instance, Refs. [169, 170]). With the aim of explaining the PTA data, in this work, we focus on the generation GWs that are associated with the formation of PBHs during the phase of reheating described by a generic EoS. We shall consider an inflationary scalar power spectrum with a peak as given by Eq. (2.8) and compare the resulting spectral density of induced GWs with the NANOGrav 15-year data. Since we are interested in understanding the compatibility of the GW background generated during reheating with the detected NANOGrav 15-year signal, we shall assume a considerably low reheating temperature of $T_{\mathrm{re}}\simeq 50\,\mathrm{MeV}$. Such an assumption ensures that the wave numbers with $k>k_{\mathrm{re}}$ fall in the NANOGrav band of frequencies. Further, it is known that, when $w_{\mathrm{re}}>1/3$, a prolonged period of reheating can cause the spectral density of primary GWs (i.e. the GWs wave generated from the vacuum during inflation) to cross the BBN bound. Later, we shall also utilize the bound to constrain the tensor-to-scalar ratio generated during inflation.
@@ 3.1 Spectral density of secondary GWs over wave numbers that re-enter during the phase of reheating
As we mentioned earlier when the primordial scalar power spectrum is boosted on small scales to increase the number of PBHs formed, the enhanced scalar power also induces secondary GWs of significant strengths. In this section, we shall discuss the evaluation of the spectral energy density of the scalar-induced, secondary GWs generated during the phase of reheating and the epoch of radiation domination. We shall closely follow the formalism developed in the earlier efforts in this context (see, for instance, Refs. [134, 135, 20]).
@@
where $x=k\,\eta$, and the coefficients $C$ and $D$ are to be determined by matching the transfer function and its time derivative with the earlier epoch. Note that, the limit $x\ll 1$ corresponds to early times. In such a limit, since the second term in the above expression for $T(x)$ diverges, we shall set $D=0$, i.e. we shall assume that there are no decaying solutions. Conventionally, the transfer function is chosen to be $T\to 1$ at large scales or, equivalently, when $x\ll 1$. Such a condition determines the value of the coefficient $C$ to be $C=(2/\sqrt{w_{\mathrm{re}}})^{\alpha+3/2}\,\Gamma(\alpha+5/2)$ so that the transfer function is given by
@@
Note that, since, in our analysis, we assume that the wave numbers near the peak in the scalar power spectrum re-enter during the phase of reheating, over the wave numbers $k_{\mathrm{re}}<k<k_{\mathrm{e}}$, the dominant contribution to the spectral density of secondary GWs can be expected to arise from the phase. The phase of reheating will be followed by the epoch of radiation domination. The secondary GWs generated over $k_{\mathrm{re}}<k<k_{\mathrm{e}}$ during reheating will freely propagate during radiation domination. In the following section, we shall discuss the effects of the generation of secondary GWs during the epoch of radiation domination.
@@ 3.2 Spectral density of secondary GWs over wave numbers that re-enter during radiation domination
The more general form of the kernel during radiation domination—which accounts for the prior phase of reheating—can be derived with $x_{\mathrm{re}}$ redefined as $x_{\mathrm{re}}/2\to[\alpha/(1+\alpha)]\,x_{\mathrm{re}}$ (in this regard, see Ref. [176], App. A). Moreover, the constants $A(x_{\mathrm{re}})$ and $B(x_{\mathrm{re}})$ in the transfer function (3.23) above can be determined by matching the Bardeen potential and its time derivative at $\eta_{\mathrm{re}}$. Clearly, the resulting kernel will reduce to the standard kernel in the epoch of radiation domination for wave numbers $k\ll k_{\mathrm{re}}$ that re-enter much later than $\eta_{\mathrm{re}}$. In the late time limit. i.e. as $x\to\infty$, we can write the kernel during radiation domination as [169, 170]
@@
The dimensionless spectral density parameter $\Omega_{{}_{\mathrm{GW}}}(k,\eta)$ associated with $\rho_{{}_{\mathrm{GW}}}(k,\eta)$ is defined with respect to the critical density at the time, viz. $\rho_{\mathrm{cr}}(\eta)$, as
@@
The present-day, dimensionless spectral density of secondary GWs can be expressed in terms of the quantity $\Omega_{{}_{\mathrm{GW}}}(k,\eta)$ as follows:
@@
where $c_{g}=\left(g_{k}/g_{0}\right)\left(g_{\mathrm{s}0}/g_{\mathrm{s}k}\right)^{4/3}$, with $(g_{k},g_{\mathrm{s}k})$ and $(g_{0},g_{\mathrm{s}0})$ denoting the effective number of relativistic degrees of freedom associated with the energy density of radiation and entropy when the wave number $k$ re-enters the Hubble radius and today, respectively.
@@
Figure 2: The dimensionless spectral energy density $\Omega_{{}_{\mathrm{GW}}}$ of both the primary (as dashed curves) and secondary (as solid curves) GWs arising in the scenario of interest are plotted as a function of the frequency $f$. In plotting the spectral density of secondary GWs, we have chosen the parameters of the primordial scalar power spectrum [cf. Eq. (2.8)] to be $\log_{10}A_{0}=-1.5$, $\log_{10}(k_{\mathrm{peak}}/\mathrm{Mpc}^{-1})=7.5$ and $n_{0}=-2$. We have set the reheating temperature to be $T_{\mathrm{re}}=50\,\mathrm{MeV}$ and have plotted the results for the following values of the EoS parameter during reheating: $w_{\mathrm{re}}=(1/10,1/3,2/5,3/5,9/10)$. We should mention that the primary spectral densities are plotted for the maximum value of the tensor-to-scalar ratio allowed by the CMB data, viz. $r=0.036$ [114], which sets the upper bound on the energy scale of inflation (in this context, see our discussion in Sec. 3.4) We have also indicated the so-called BBN bound (in dashed red line), which limits the total energy density of relativistic fields during the epoch of radiation domination.
@@
In the figure, we have also included the projected sensitivity curves of different current and forthcoming GW observatories such as PTA, Square Kilometre Array (SKA) [177], LISA [178], DECIGO [179, 180], BBO [181, 182, 183], MAGIS-100 [169, 184], advanced LIGO + Virgo [185, 186], ET [187], and CE [188]. We should point out that the sensitivity curves have been arrived at by assuming a power law for the spectral density of GWs and integrating the noise data over a range of frequencies and observation time. The sensitivity curves indicate in a simple way whether the spectral densities we have calculated will be detected by the GW observatories. Note that, in plotting the dimensionless spectral densities of the primary and secondary GWs, we have fixed the reheating temperature to be $T_{\mathrm{re}}\simeq 50\,\mathrm{MeV}$. Moreover, in the case of the secondary GWs, we have assumed that the scalar power spectrum has a broken power law form, as given in Eq. (2.8). Further, we have chosen that the peak in the scalar power spectrum to be at $k_{\mathrm{peak}}=10^{7.5}\,\mathrm{Mpc}^{-1}$, and we have assumed that $n_{0}=-2$. The reason for choosing these values of $T_{\mathrm{re}}$ and $k_{\mathrm{peak}}$ is to effectively constrain $w_{\mathrm{re}}$ with the aid of the PTA data. Recall that, we have assumed a sudden transition from the phase of reheating to the era of radiation domination. We have chosen the above-mentioned parameters such that the wave number $k_{\mathrm{re}}$ occurs before the range of frequencies that the PTAs are sensitive to. In the figure, we have plotted the quantity $\Omega_{{}_{\mathrm{GW}}}(f)$ for the following five values of the parameter that describes the EoS during reheating: $w_{\mathrm{re}}=(1/10,1/3,2/5,3/5,9/10)$. It is clear from the figure that, as $w_{\mathrm{re}}$ increases, the height of the peak in the secondary $\Omega_{{}_{\mathrm{GW}}}(f)$ also increases and the bump-like feature around the peak is flattened. It is also evident from the figure that, for all values of $w_{\mathrm{re}}$, prior to the peak, the spectral density of the secondary GWs behaves as expected for wave numbers that re-enter the Hubble radius during the epoch of radiation domination. Beyond the peak, the spectral density of the secondary GWs decreases in different manner for different values of $w_{\mathrm{re}}$. As $w_{\mathrm{re}}$ is increased from zero, $\Omega_{{}_{\mathrm{GW}}}(f)$ falls less and less and it becomes nearly constant for $w_{\mathrm{re}}=1/3$, and starts increasing again for $w_{\mathrm{re}}>1/3$. This happens due to the relative redshift between the background and GWs. The second rise in the spectral density of secondary GWs can also be detected by observatories such as LISA, DECIGO, and BBO. Hence, the presence of a double peak can also indicate an EoS with $w_{\mathrm{re}}>1/3$.
@@ 3.3 Comparison with the NANOGrav 15-year data
In this section, we scan our parameter space to find the region of the space that optimally fits the PTA data. We further examine whether the optimal values for the parameters we obtain are consistent with the available constraints on $f_{{}_{\mathrm{PBH}}}(M)$.
@@
Refer to captionRefer to caption
Figure 3: We have plotted the marginalized posterior distributions of the parameters that have been arrived at upon comparing the models R4pF (on the left) and R3pF (on the right) with the NANOGrav 15-year data. Note that, in both the models, while the data constrains the parameters $A_{0}$ and $k_{\mathrm{peak}}$ fairly well (as is suggested by the peaks in their posterior distributions), the constraints on the parameters $w_{\mathrm{re}}$ and $n_{0}$ (in the case of the model R4pF) are rather weak.
@@
Table 2: We have listed the Bayesian factors $\mathrm{BF}_{Y,X}$ we have obtained for the model R2pB that invokes primordial physics as the source of the stochastic GW background observed by the NANOGrav 15-year data, when compared to the astrophysical scenario of SMBHBs. Bayesian factors $\mathrm{BF}_{Y,X}$ that far exceed unity indicate strong evidence for the model $Y$ with respect to the model $X$. Clearly, when $\delta_{\mathrm{c}}=\delta_{\mathrm{c}}^{\mathrm{an}}$ and $\delta_{\mathrm{c}}=1.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$, the NANOGrav 15-year data strongly favors the model R2pB when compared to the model of SMBHBs.
@@
Figure 5: The spectral density of secondary GWs $\Omega_{{}_{\mathrm{GW}}}(f)$ that correspond to the best-fit values of the parameters in the four models, viz. R4pF, R3pF, R3pB and R2pB, are plotted. In the figure, we have also included the probability densities from the Kernel Density Estimators (KDEs) of the NANOGrav 15-year data (as gray violins) [194]. It is visually evident that, when $\delta_{\mathrm{c}}=0.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$, the spectral densities corresponding to the latter two models do not fit the data well.
@@
Refer to captionRefer to caption
Figure 6: The inflationary scalar power spectrum $\mathcal{P}_{\mathcal{R}}(k)$ (on the left) and the fraction of PBHs that constitute the energy density of cold dark matter today $f_{{}_{\mathrm{PBH}}}(M)$ (on the right), which correspond to the best-fit values of the parameters in the four models are plotted. Clearly, in the case of the models R4pF and R3pF, the best-fit values lead to the overproduction of PBHs. It is also interesting to note that the best-fit inflationary scalar power spectra avoid the current/expected bounds from spectral distortions.

### v2

@@ 3 Generation of scalar-induced secondary GWs and the PTA data
As is well known, primordial GWs are the transverse and traceless components of the tensor perturbations in the metric describing the FLRW universe. There can be different origins for GWs in the early universe. At the linear order in perturbation theory, the scalar and tensor perturbations evolve independently. However, the tensor perturbations at the second order can be sourced by the scalar perturbations at the first order. As a result, whenever the amplitudes of the scalar perturbations are enhanced on small scales so that the PBHs are formed abundantly, they induce tensor perturbations of significant strengths at the second order, often leading to a peak in the spectral density of GWs (see, for instance, Refs. [196, 197]). With the aim of explaining the PTA data, in this work, we focus on the generation GWs that are associated with the formation of PBHs during the phase of reheating described by a generic EoS. We shall consider an inflationary scalar power spectrum with a peak as given by Eq. (2.8) and compare the resulting spectral energy density of induced GWs with the NANOGrav 15-year data. Since we are interested in understanding the compatibility of the GW background generated during reheating with the detected NANOGrav 15-year signal, we shall assume a considerably low reheating temperature of $T_{\mathrm{re}}\simeq 50\,\mathrm{MeV}$. Such an assumption ensures that the wave numbers with $k>k_{\mathrm{re}}$ fall in the NANOGrav band of frequencies. Further, it is known that, when $w_{\mathrm{re}}>1/3$, a prolonged period of reheating can cause the spectral energy density of primary GWs (i.e. the GWs wave generated from the vacuum during inflation) to cross the so-called BBN bound. Later, we shall also utilize the bound to constrain the tensor-to-scalar ratio generated during inflation.
@@ 3.1 Spectral density of secondary GWs over wave numbers that re-enter during the phase of reheating
As we mentioned earlier, when the primordial scalar power spectrum is boosted on small scales to increase the number of PBHs formed, the enhanced scalar power also induces secondary GWs of significant strengths. In this section, we shall discuss the evaluation of the spectral energy density of the scalar-induced, secondary GWs generated during the phase of reheating and the epoch of radiation domination. We shall closely follow the formalism developed in the earlier efforts in this context (see, for instance, Refs. [139, 140, 20]).
@@
where $x=k\,\eta$, and the coefficients $C$ and $D$ are to be determined by matching the transfer function and its time derivative with the earlier epoch. Note that the limit $x\ll 1$ corresponds to early times. In such a limit, since the second term in the above expression for $T(x)$ diverges, we shall set $D=0$, i.e. we shall assume that there are no decaying solutions. Conventionally, the transfer function is chosen to be $T\to 1$ at large scales or, equivalently, when $x\ll 1$. Such a condition determines the value of the coefficient $C$ to be $C=(2/\sqrt{w_{\mathrm{re}}})^{\alpha+3/2}\,\Gamma(\alpha+5/2)$ so that the transfer function is given by
@@
Note that, since, in our analysis, we assume that the wave numbers near the peak in the scalar power spectrum re-enter during the phase of reheating (i.e. they lie in the range $k_{\mathrm{re}}<k<k_{\mathrm{e}}$), the dominant contribution to the spectral density of secondary GWs can be expected to arise from the phase. The phase of reheating will be followed by the epoch of radiation domination. The secondary GWs generated over $k_{\mathrm{re}}<k<k_{\mathrm{e}}$ during reheating will freely propagate during radiation domination. In the following section, we shall discuss the generation of secondary GWs during the epoch of radiation domination.
@@ 3.2 Spectral density of secondary GWs over wave numbers that re-enter during radiation domination
The more general form of the kernel during radiation domination—which accounts for the prior phase of reheating—can be derived with $x_{\mathrm{re}}$ redefined as $x_{\mathrm{re}}/2\to[\alpha/(1+\alpha)]\,x_{\mathrm{re}}$ (in this regard, see Ref. [203], App. A). Moreover, the constants $A(x_{\mathrm{re}})$ and $B(x_{\mathrm{re}})$ in the transfer function (3.23) above can be determined by matching the Bardeen potential and its time derivative at $\eta_{\mathrm{re}}$. Clearly, the resulting kernel will reduce to the standard kernel in the epoch of radiation domination for wave numbers $k\ll k_{\mathrm{re}}$ that re-enter much later than $\eta_{\mathrm{re}}$. In the late time limit, i.e. as $x\to\infty$, we can write the kernel during radiation domination as [196, 197]
@@
The dimensionless spectral energy density parameter $\Omega_{{}_{\mathrm{GW}}}(k,\eta)$ associated with $\rho_{{}_{\mathrm{GW}}}(k,\eta)$ is defined with respect to the critical density at the time, viz. $\rho_{\mathrm{cr}}(\eta)$, as
@@
The present-day, dimensionless spectral energy density of secondary GWs can be expressed in terms of the quantity $\Omega_{{}_{\mathrm{GW}}}(k,\eta)$ as follows:
@@
where $c_{g}=\left(g_{k}/g_{0}\right)\left(g_{\mathrm{s,}0}/g_{\mathrm{s},k}\right)^{4/3}$, with $(g_{k},g_{\mathrm{s},k})$ and $(g_{0},g_{\mathrm{s},0})$ denoting the effective number of relativistic degrees of freedom associated with the energy density of radiation and entropy when the wave number $k$ re-enters the Hubble radius and today, respectively.
@@
Figure 2: The dimensionless spectral energy density $\Omega_{{}_{\mathrm{GW}}}$ of both the primary (as dashed curves) and secondary (as solid curves) GWs arising in the scenario of interest are plotted as a function of the frequency $f$. We have plotted the spectral energy densities up to the frequency corresponding to the wavenumber $k_{\mathrm{e}}$ which leaves the Hubble radius at the end of inflation. In plotting the spectral density of secondary GWs, we have chosen the parameters of the primordial scalar power spectrum [cf. Eq. (2.8)] to be $\log_{10}A_{0}=-1.5$, $\log_{10}(k_{\mathrm{peak}}/\mathrm{Mpc}^{-1})=7.5$ and $n_{0}=-2$. We have set the reheating temperature to be $T_{\mathrm{re}}=50\,\mathrm{MeV}$ and have plotted the results for the following values of the EoS parameter during reheating: $w_{\mathrm{re}}=(1/10,1/3,2/5,3/5,9/10)$. We should mention that the spectral energy densities of the primary GWs are plotted corresponding to the maximum value of the tensor-to-scalar ratio allowed by the CMB data, viz. $r=0.036$ [117], which sets the upper bound on the energy scale of inflation (in this context, see our discussion in Sec. 3.4). In the figure, we have also indicated the so-called BBN bound (as the dashed red line), which limits the total energy density of relativistic fields during the epoch of radiation domination.
@@
In the figure, we have also included the projected sensitivity curves of different current and forthcoming GW observatories such as PTA, Square Kilometre Array (SKA) [204], LISA [205], DECIGO [206, 207], BBO [208, 209, 210], MAGIS-100 [196, 211], advanced LIGO + Virgo (LVK) [212, 213], ET [214], and CE [215]. We should point out that the sensitivity curves have been arrived at by assuming a power law for the spectral density of GWs and integrating the noise data over a range of frequencies and observation time. The sensitivity curves indicate in a simple way whether the spectral densities we have calculated will be detected by the GW observatories. Note that, in plotting the dimensionless spectral energy densities of the primary and secondary GWs, we have fixed the reheating temperature to be $T_{\mathrm{re}}\simeq 50\,\mathrm{MeV}$. Moreover, in the case of the secondary GWs, we have assumed that the scalar power spectrum has a broken power law form, as given in Eq. (2.8). Further, we have chosen that the peak in the scalar power spectrum to be at $k_{\mathrm{peak}}=10^{7.5}\,\mathrm{Mpc}^{-1}$, and we have assumed that $n_{0}=-2$. The reason for choosing these values of $T_{\mathrm{re}}$ and $k_{\mathrm{peak}}$ is to effectively constrain $w_{\mathrm{re}}$ with the aid of the PTA data. Recall that, we have assumed a sudden transition from the phase of reheating to the era of radiation domination. We have chosen the above-mentioned parameters such that the wave number $k_{\mathrm{re}}$ occurs before the range of frequencies that the PTAs are sensitive to. In the figure, we have plotted the quantity $\Omega_{{}_{\mathrm{GW}}}(f)$ for the following five values of the parameter that describes the EoS during reheating: $w_{\mathrm{re}}=(1/10,1/3,2/5,3/5,9/10)$. It is clear from the figure that, as $w_{\mathrm{re}}$ increases, the height of the peak in the secondary $\Omega_{{}_{\mathrm{GW}}}(f)$ also increases and the bump-like feature around the peak is flattened. It is also evident from the figure that, for all values of $w_{\mathrm{re}}$, prior to the peak, the spectral density of the secondary GWs behaves as expected for wave numbers that re-enter the Hubble radius during the epoch of radiation domination. Beyond the peak, the spectral density of the secondary GWs decreases in different manner for different values of $w_{\mathrm{re}}$. As $w_{\mathrm{re}}$ is increased from zero, $\Omega_{{}_{\mathrm{GW}}}(f)$ falls less and less and it becomes nearly constant for $w_{\mathrm{re}}=1/3$, and starts increasing again for $w_{\mathrm{re}}>1/3$. This happens due to the relative redshift between the background and GWs. The second rise in the spectral density of secondary GWs can also be detected by observatories such as LISA, DECIGO, and BBO. Hence, the presence of a double peak can indicate an EoS with $w_{\mathrm{re}}>1/3$.
There are a couple of related points regarding Fig. 2 that require further emphasis. In the figure, we have plotted the spectral energy densities of the primary and secondary GWs up to the frequency corresponding to the wavenumber $k_{\mathrm{e}}$, where, recall that, $k_{\mathrm{e}}$ is the wavenumber that leaves the Hubble radius at the end of inflation. Note that the Fourier modes with wavenumbers larger than $k_{\mathrm{e}}$ always stay within the Hubble radius and hence they do not contribute to the energy density of GWs (for a discussion in this regard, see, for example, Refs. [216, 217]). It is clear from the figure that, for $w_{\mathrm{re}}=9/10$, the spectral energy density of primary GWs crosses the sensitivity curve of LVK as well as the BBN bound from $\Delta N_{\mathrm{eff}}$. Whereas, for $w_{\mathrm{re}}=3/5$, the spectral energy density of the primary GWs crosses only the BBN bound. Based on the fact that LVK did not observe a SGWB [22], we can exclude the spectral energy densities passing through it. Therefore, for the tensor-to-scalar ratio of $r=0.036$, we can rule out the possibility of a prolonged reheating phase with either $w_{\mathrm{re}}=9/10$ or $w_{\mathrm{re}}=3/5$. But, a smaller value of $r$ (or, equivalently, a lower inflationary energy scale) would still remain compatible with such reheating histories. These suggest that, for a given energy scale of inflation, BBN and LVK can also be used to constrain the parameters that characterize the epoch of reheating. In fact, blue-tilted inflationary spectra can be constrained by LVK very effectively. However, we should point out that, if we consider a flat or red-tilted inflationary tensor power spectrum, the bound from $\Delta N_{\mathrm{eff}}$ provides a stronger constraint on primary GWs than LVK.
@@ 3.3 Comparison with the NANOGrav 15-year data
In this section, we scan our parameter space to find the region of the space that optimally fits the PTA data. We shall further examine whether the optimal values for the parameters we obtain are consistent with the available constraints on $f_{{}_{\mathrm{PBH}}}(M)$.
@@
Figure 3: We have plotted the marginalized posterior distributions of the parameters that have been arrived at upon comparing the models R4pF (on the left) and R3pF (on the right) with the NANOGrav 15-year data. Note that, in both the models, while the data constrains the parameters $A_{0}$ and $k_{\mathrm{peak}}$ fairly well (as is suggested by the peaks in their posterior distributions), the constraints on the parameters $w_{\mathrm{re}}$ and $n_{0}$ (in the case of the model R4pF) are rather weak. In fact, in the case of $n_{0}$, the posterior distribution is flat over the range of priors that we have worked with [cf. Tab. 3.3].
@@
Table 2: We have listed the Bayesian factors $\mathrm{BF}_{Y,X}$ that we have obtained for the model R2pB which invokes primordial physics as the source of the SGWB observed by the NANOGrav 15-year data, when compared to the astrophysical scenario of SMBHBs. Bayesian factors $\mathrm{BF}_{Y,X}$ that far exceed unity indicate strong evidence for the model $Y$ with respect to the model $X$. Clearly, when $\delta_{\mathrm{c}}=\delta_{\mathrm{c}}^{\mathrm{an}}$ and $\delta_{\mathrm{c}}=1.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$, the NANOGrav 15-year data strongly favors the model R2pB when compared to the model of SMBHBs.
@@
Figure 5: The spectral energy density of secondary GWs $\Omega_{{}_{\mathrm{GW}}}(f)$ that correspond to the best-fit values of the parameters in the four models, viz. R4pF, R3pF, R3pB and R2pB, are plotted. In the figure, we have also included the probability densities from the Kernel Density Estimators (KDEs) of the NANOGrav 15-year data (as gray violins) [223]. It is visually evident that, when $\delta_{\mathrm{c}}=0.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$, the spectral energy densities corresponding to the latter two models do not fit the data well.
@@
Figure 6: The inflationary scalar power spectrum $\mathcal{P}_{\mathcal{R}}(k)$ (on the left) and the fraction of PBHs that constitute the energy density of cold dark matter today $f_{{}_{\mathrm{PBH}}}(M)$ (on the right), which correspond to the best-fit values of the parameters in the four models are plotted. Clearly, in the case of the models R4pF and R3pF, the best-fit values lead to the overproduction of PBHs. It is also interesting to note that the best-fit inflationary scalar power spectra avoid the current and expected bounds from spectral distortions.